# From Strangers to Assistants: Fast Desire Alignment for Embodied Agent-User Adaptation

### v1

@@ Abstract
While embodied agents have made significant progress in performing complex physical tasks, real-world applications demand more than pure task execution. The agents must collaborate with unfamiliar agents and human users, whose goals are often vague and implicit. In such settings, interpreting ambiguous instructions and uncovering underlying desires is essential for effective assistance. Therefore, fast and accurate desire alignment becomes a critical capability for embodied agents. In this work, we first develop a home assistance simulation environment HA-Desire that integrates an LLM-driven human user agent exhibiting realistic value-driven goal selection and communication. The ego agent must interact with this proxy user to infer and adapt to the user’s latent desires. To achieve this, we present a novel framework FAMER for fast desire alignment, which introduces a desire-based mental reasoning mechanism to identify user intent and filter desire-irrelevant actions. We further design a reflection-based communication module that reduces redundant inquiries, and incorporate goal-relevant information extraction with memory persistence to improve information reuse and reduce unnecessary exploration. Extensive experiments demonstrate that our framework significantly enhances both task execution and communication efficiency, enabling embodied agents to quickly adapt to user-specific desires in complex embodied environments.
@@ 1 Introduction
Embodied AI has seen rapid progress in recent years, driven by the collection of datasets brohan2022rt ; brohan2023rt ; o2024open ; fang2023rh20t ; wang2025adamanip and the development of large vision-language-action (VLA) models black2024pi_0 ; driess2023palm ; kim2024openvla ; team2024octo . These advances have paved the way for general-purpose robots capable of performing complex manipulation tasks in the physical world. However, real-world deployment of embodied agents requires more than physical capabilities. It also demands the ability to interact effectively with diverse human users.
@@
A prime example is the home assistant robots. Even if these robots are trained on broad human-centric datasets, they inevitably face unfamiliar users whose specific values and preferences are unknown at deployment. To offer effective assistance, the agent must infer and adapt to these user-specific attributes. In this way, the agent minimizes repetitive communication and demonstrates proactive, personalized behavior, similar to a considerate human assistant. For instance, as illustrated in Figure 1, a robot enters a new home without prior knowledge of the resident. Over time, it gradually learns that the user is allergic to caffeine and prefers something refreshing for breakfast due to poor sleep caused by a heavy workload. Therefore, the robot infers that the user wants juice and serves it without needing to ask. Such behavior highlights the necessity of rapid and accurate desire alignment in embodied assistance, enabling robots to build trust and deliver truly helpful service.
@@
Previous works have investigated collaboration with unfamiliar partners under the paradigms of ad-hoc teamwork (AHT) rahman2021towards ; ravula2019ad ; stone2010ad and zero-shot coordination (ZSC) cui2021k ; hu2021off ; hu2020other . However, these efforts have primarily focused on simplified domains such as board games like Hanabi bard2020hanabi and 2D grid-based simulations like Level-Based Foraging albrecht2015game and Overcooked carroll2019utility . These environments lack human-like, value-driven goal specification, natural communication and embodied actions, limiting their applicability to realistic embodied agent-user adaptation scenarios.
@@
To bridge this gap, we first introduce a new embodied simulation environment, HA-Desire (Home Assistance with diverse Desire), which is built upon the VirtualHome platform puig2018virtualhome . HA-Desire features rich 3D household scenes and diverse objects, supporting various tasks such as preparing an afternoon snack or setting the dinner table. Crucially, it includes an LLM-driven proxy human user that samples goals based on randomly assigned value attributes and vague task descriptions. The user interacts with the ego agent via natural language and does not explicitly state its goals. Instead, it offers indirect, human-like hints such as “I want something sweet and crunchy.” This setup requires the ego agent to perform strategic inference and interactive reasoning to uncover the user’s latent desires and execute a task that fully satisfies them.
@@
To tackle this challenging problem of fast agent-user adaptation, we propose FAMER (Fast Adaptation via MEntal Reasoning), a novel framework that leverages the reasoning capabilities and commonsense knowledge of large vision-language models to improve both communication efficiency and task execution. At the core of FAMER is a desire-centered mental reasoning module that extracts confirmed goals from user messages and infers the user’s underlying mental state, including values, preferences, and latent desires. To reduce redundant communications, FAMER also incorporates a reflection-based communication mechanism that prompts the agent to reason about what has already been learned and to ask only for missing or unconfirmed information. Additionally, FAMER includes a goal-relevant information extraction module that identifies critical task-related details, such as object containers or room locations, and stores them in a persistent memory across episodes. This enables the agent to reuse previously gathered knowledge and avoid unnecessary exploration. Together, these components allow the ego agent to rapidly align with the user’s desires and plan efficiently in complex, multi-step embodied tasks.
@@
We evaluate FAMER on two representative tasks: Snack & Table, in our HA-Desire environment. Each task is tested under two settings: Medium and Large, denoting the number of goals to satisfy. Extensive quantitative and qualitative experiments show that FAMER significantly outperforms baselines in task completion score, execution efficiency, and communication cost. Ablation studies further highlight the contribution of each key component in the framework.
@@
  • We formulate the novel problem of rapid adaptation to value-driven, unknown users in embodied settings, and introduce HA-Desire, a 3D simulation environment featuring naturalistic user interactions for evaluating agent-user adaptation.
@@
Figure 1: An illustration of Embodied Agent-User Adaptation. The embodied home assistant robot encounters a new user with unknown values and preferences. Through interaction over time, the agent learns the user’s aversion to caffeine and preference for refreshing drinks in the morning due to inadequate sleep. By aligning with the user’s implicit desires, the agent proactively serves juice without being explicitly instructed, demonstrating high-quality assistant service.
@@ 2 Related Work
Value Alignment Value alignment has been extensively studied in both language modeling and agent design. In the context of LLMs, alignment techniques such as RLHF ouyang2022training ; dai2023safe ; ji2023ai aim to align model outputs with human preferences, but these efforts primarily focus on static, text-based tasks and do not address the challenges of dynamic, embodied interactions. In human-AI collaboration, value alignment involves inferring user preferences through observation or feedback yuan2022situ ; hiatt2017human ; fisac2020pragmatic . More closely related to our setting are mental reasoning agents inspired by Theory of Mind rabinowitz2018machine ; wang2021tomc , which model other agents’ beliefs and desires to support assistance. D2A wang2024simulating simulates human desires using LLMs, but is limited to text-based environments. CHAIC du2024constrained introduces an embodied social intelligence challenge that focuses on reasoning under physical constraints, but does not address the diversity of human values and goals. In contrast, our work introduces an integrated embodied simulation platform with naturalistic, value-driven goal generation and communication.
Adaptive Agents Adaptation in multi-agent settings has been studied under the paradigms of zero-shot coordination (ZSC) hu2020other ; hu2021off ; cui2021k ; lupu2021trajectory ; strouse2021collaborating and ad-hoc teamwork (AHT) stone2010ad ; rahman2021towards ; chen2020aateam ; mirsky2020penny ; ma2024fast , where agents must coordinate with unseen partners without prior agreement. While these approaches are effective in structured domains such as Hanabi bard2020hanabi and Overcooked carroll2019utility , they rely on symbolic observations, making them less suitable for complex, embodied human-agent collaboration. More recently, LLM-based agent frameworks li2023camel ; yao2023react ; zhang2023building ; zhang2024proagent ; liu2024capo have demonstrated impressive capabilities in reasoning and planning within interactive environments. However, most assume known goals or static user preferences and lack mechanisms for inferring latent values through interaction. Our work builds on this line by addressing the challenge of fast adaptation to unknown, value-driven user goals via desire inference, memory utilization, and efficient, human-like dialogue in rich embodied tasks.
@@ 3 Embodied Home Assistance Simulation Environment
We present HA-Desire (Home Assistance with diverse Desire), a novel embodied simulation environment designed to study user-specific adaptation in realistic home scenarios. As illustrated in Figure 2, HA-Desire is built upon a richly detailed 3D environment populated with diverse object assets and room layouts, allowing for visually grounded tasks with high variability.
To simulate real-world agent-user adaptation challenges, we instantiate a proxy human user within the environment. This user is endowed with a set of latent value attributes and receives only a vague task description. Based on these values, the user samples a set of desire-related goals from a potential goal set. Importantly, the user does not directly reveal these goals but instead provides indirect hints about preferences and intentions in response to the ego agent’s inquiries.
@@ 3.1 Problem Formulation
In the desire-centered agent-user adaptation problem, the task goal is not explicitly fixed, but the number of goals is predefined. Instead, there exists a potential goal set $G_{p}$, from which the task goal must be sampled by the human user $H$ and inferred by the ego agent $E$. Given a vague task description $T$, the proxy human user first samples a set of value attributes $V$ from a predefined value space. Based on these values and the task description, the user then samples a set of desire-driven goals $G=G(V,G_{p},T)\subset G_{p}$.
@@ 3.2 Value-driven Human User
As illustrated in Figure 2, the proxy human user begins by sampling discretized value attributes $V$ from a predefined task-related value space. For example, in the Prepare Snack task, the value space spans five dimensions: Hungry, Thirsty, SweetTooth, Fruitarian, and Alcoholic, each taking on one of three discrete levels—Not, Somewhat, or Very. These attributes reflect user’s latent desires.
@@
To simulate natural and indirect human communication, the user is constrained via LLM prompting not to reveal the true goal set directly. Instead, it responds with value-driven hints that reflect its preferences and desires. This encourages the ego agent to reason about the user’s intent through interaction, rather than relying on explicit instructions. The detailed LLM prompting strategy used to sample goals and generate user responses is provided in the Appendix.
Figure 2: Overview of the HA-Desire environment. The simulation environment contains diverse objects and scenes. The proxy human user samples value attributes from a predefined space, which guides goal selection from a potential goal set via LLM. The user is constrained from directly revealing the true goals and instead communicates through desire-centered hints. This setup encourages the ego agent to infer user intent through interactive reasoning rather than relying on explicit instructions.
Figure 3: Overview of the FAMER framework. FAMER comprises three key components: KeyInfo Extraction, Desire-Centered Mental Reasoning (including Goal Confirmation and Desire Inference), and Efficient Communication. These are supported by a memory module, perception module, and planning module, which together form an integrated pipeline for embodied agent-user adaptation.
4 Method
To enable fast and effective adaptation to user-specific desires, we propose a novel framework, FAMER (Fast Adaptation via MEntal Reasoning), which leverages the reasoning capabilities and commonsense knowledge embedded in large VLMs. In this work, we utilize GPT-4o hurst2024gpt .
As illustrated in Figure 3, FAMER integrates three core components: Key Information Extraction, Desire-Centered Mental Reasoning, and Efficient Communication. These modules work in concert to help the ego agent infer user intent, plan accurately, and minimize redundant interactions. We detail each component in the following subsections.
@@ 4.1 Information Extraction
In HA-Desire, the ego agent receives first-person RGB-D images as observations. To extract structured information from these inputs, we employ a perception module based on Mask R-CNN he2017mask , trained on collected scene images zhang2023building . The module first predicts instance segmentation masks from the RGB image and then constructs 3D point clouds using the RGB-D data. These outputs are used to build a scene graph that encodes object locations, categories, and spatial relationships.
@@ 4.2 Desire-centered Mental Reasoning
The Goal Confirmation component extracts confirmed goals from the user’s responses by VLM reasoning. For example, if the agent asks, “Is juice what you want?” and the user replies, “Correct! Try to look for something crunchy,” the system confirms that juice is one of the user’s desired items. This process grounds part of the goal set and reduces future uncertainty.
@@ 4.3 Efficient Communication
When communication is necessary, the agent formulates targeted, desire-aligned questions aimed at resolving specific ambiguities. For example, if the agent has inferred a preference for sweet items but is uncertain about texture, it may ask, “Is cupcake what you want?” instead of issuing vague or open-ended queries. This focused interaction strategy minimizes the communication burden on the user while enabling the agent to efficiently acquire high-value information.
@@ 5.1 Tasks & Metrics
We evaluate FAMER in two representative tasks instantiated within our proposed HA-Desire environment: Prepare Afternoon Snack and Set Up Dinner Table. Each task is associated with a five-dimensional value space that governs user preferences. The Snack task includes 10 potential goals, while the Table task contains 8. Each task is further divided into two levels: Medium and Large. In Medium setting, the number of target goals is 2, and the maximum episode length is 200 steps. In Large setting, the agent must satisfy 4 goals within 300 steps. For example, the Snack-M task involves a total of $C(10,2)=45$ possible goal combinations. The agent is allowed to interact with the user for 3 episodes, and the user’s value attributes remain fixed throughout the entire interaction.
@@
Score: Given $N$ total goals, the agent receives a reward of $\frac{1}{N}$ for each correct goal achieved. Completing an incorrect or distracting goal incurs a penalty of $-\frac{1}{2N}$. The maximum achievable score per episode is 1, corresponding to the completion of all goals without any mistakes.
Step: The total number of environment steps taken by the agent within a single episode.
@@
Notably, HA-Desire supports diverse objects and scenes and goal generation is automated, making it easy to extend to other tasks. Further details of the tasks are provided in the Appendix.

### v2

# Communication-Efficient Desire Alignment for Embodied Agent-Human Adaptation

@@ Abstract
While embodied agents have made significant progress in performing complex physical tasks, real-world applications demand more than pure task execution. The agents must collaborate with unfamiliar agents and human users, whose goals are often vague and implicit. In such settings, interpreting ambiguous instructions and uncovering underlying desires is essential for effective assistance. Therefore, fast and accurate desire alignment becomes a critical capability for embodied agents. In this work, we first develop a home assistance simulation environment HA-Desire that integrates an LLM-driven proxy human user exhibiting realistic value-driven goal selection and communication. The ego agent must interact with this proxy user to infer and adapt to the user’s latent desires. To achieve this, we present a novel framework FAMER for fast desire alignment, which introduces a desire-based mental reasoning mechanism to identify user intent and filter desire-irrelevant actions. We further design a reflection-based communication module that reduces redundant inquiries, and incorporate goal-relevant information extraction with memory persistence to improve information reuse and reduce unnecessary exploration. Extensive experiments demonstrate that our framework significantly enhances both task execution and communication efficiency, enabling embodied agents to quickly adapt to user-specific desires in complex embodied environments.
@@ 1 Introduction
Embodied AI has seen rapid progress in recent years, driven by the collection of large datasets Brohan et al. (2022; 2023); O’Neill et al. (2024); Fang et al. (2023); Wang et al. (2025b) and the development of large vision-language-action (VLA) models Black et al. (2024); Driess et al. (2023); Kim et al. (2024); Team et al. (2024). These advances have paved the way for general-purpose robots capable of performing complex manipulation tasks in the physical world. However, real-world deployment of embodied agents requires more than physical capabilities. It also demands the ability to interact effectively with diverse human users.
@@
A prime example is the home assistant robots. Even if these robots are trained on broad human-centric datasets, they inevitably face unfamiliar users whose specific values and preferences are unknown at deployment. To offer effective assistance, the agent must infer and adapt to these user-specific attributes. In this way, the agent minimizes repetitive communication and demonstrates proactive, personalized behavior, similar to a considerate human assistant. For instance, as illustrated in Figure 1, a robot enters a new home without prior knowledge of the user. Over time, it gradually learns that the user is allergic to caffeine and prefers something refreshing for breakfast due to poor sleep caused by a heavy workload. Therefore, the robot infers that the user wants juice and serves it without needing to ask. Such behavior highlights the necessity of rapid and accurate desire alignment in embodied assistance, enabling robots to build trust and deliver truly helpful service.
@@
Previous works have investigated collaboration with unfamiliar partners under the paradigms of ad-hoc teamwork (AHT) Rahman et al. (2021); Ravula (2019); Stone et al. (2010) and zero-shot coordination (ZSC) Cui et al. (2021); Hu et al. (2021; 2020). However, these efforts have primarily focused on simplified domains such as board games like Hanabi Bard et al. (2020) and 2D grid-based simulations like Level-Based Foraging Albrecht & Ramamoorthy (2015) and Overcooked Carroll et al. (2019). These environments lack human-like, value-driven goal specification, natural communication and embodied actions, limiting their applicability to realistic embodied agent-human adaptation.
@@
To bridge this gap, we propose HA-Desire (Home Assistance with diverse Desire), a new embodied simulation environment built on VirtualHome Puig et al. (2018) that offers rich 3D household scenes, diverse objects, and tasks such as preparing an afternoon snack. Crucially, it includes an LLM-driven proxy human user that samples goals from assigned value attributes. Real human users may provide imprecise instructions, either because their goals are still being formed or because they wish to minimize communication effort. To reflect this setting, the proxy user communicates with the agent in natural language without explicitly revealing its goals, instead offering indirect hints such as “I want something sweet and crunchy.” This design forces the agent to perform strategic inference and interactive reasoning to uncover latent desires and complete tasks that satisfy the user.
@@
To tackle this challenging problem of fast agent-human adaptation, we propose FAMER (Fast Adaptation via MEntal Reasoning), a novel framework that leverages the reasoning capabilities and commonsense knowledge of large vision-language models to improve both communication efficiency and task execution. At the core of FAMER is a desire-centered mental reasoning module that extracts confirmed goals from user messages and infers the user’s underlying mental state, including values, preferences, and latent desires. To reduce redundant communications, FAMER also incorporates a reflection-based communication mechanism that prompts the agent to reason about what has already been inferred and to ask only for missing or unconfirmed information. Additionally, FAMER includes a goal-relevant information extraction module that identifies critical task-related details, such as object containers or room locations, and stores them in a persistent memory across episodes. This enables the agent to reuse previously gathered information and avoid unnecessary exploration. Together, these components allow the ego agent to rapidly align with the user’s desires and plan efficiently in complex, multi-step embodied tasks.
@@
We evaluate FAMER on two representative tasks: Snack & Table, in our HA-Desire environment. Each task is tested under two settings: Medium and Large, denoting the number of goals to satisfy. Extensive experiments with the LLM-driven proxy user and real human users show that FAMER significantly outperforms baselines in task completion score and communication efficiency. Ablation studies further highlight the contribution of each key component in the framework.
Figure 1: An illustration of Embodied Agent-Human Adaptation. The embodied home assistant robot encounters a new human user with unknown values and preferences. Through interaction over time, the agent learns the user’s aversion to caffeine and preference for refreshing drinks in the morning due to inadequate sleep. By aligning with the user’s implicit desires, the agent proactively serves juice without being explicitly instructed, demonstrating high-quality assistant service.
@@
  • We formulate the novel problem of rapid adaptation to value-driven, unknown users in embodied settings, and introduce HA-Desire, a 3D simulation environment featuring naturalistic user interactions for evaluating agent-human adaptation.
@@ 2 Related Work
Value Alignment has been extensively studied in both language models and agent design. In the context of LLMs, alignment techniques such as RLHF Ouyang et al. (2022); Dai et al. (2023); Ji et al. (2023) aim to align models with human preferences, but these efforts primarily focus on static, text-based tasks and do not address the challenges of dynamic, embodied interactions. In human-AI collaboration, value alignment involves inferring user preferences through feedback Yuan et al. (2022); Hiatt et al. (2017); Fisac et al. (2020). More closely related to our setting are mental reasoning agents inspired by Theory of Mind Rabinowitz et al. (2018); Wang et al. (2022), which model other agents’ beliefs and desires to support assistance. D2A Wang et al. (2025a) simulates human desires using LLMs, but is limited to text-based tasks. CHAIC Du et al. (2024) introduces an embodied social intelligence challenge that focuses on reasoning under physical constraints, but does not address the diversity of human values and goals. In contrast, our work introduces an embodied simulation platform with naturalistic, value-driven goal generation and communication.
Adaptive Agents. Adaptation in multi-agent settings has been studied under the paradigms of zero-shot coordination (ZSC) Hu et al. (2020; 2021); Cui et al. (2021); Lupu et al. (2021); Strouse et al. (2021) and ad-hoc teamwork (AHT) Stone et al. (2010); Rahman et al. (2021); Chen et al. (2020); Mirsky et al. (2020); Ma et al. (2024), where agents must coordinate with unseen partners without prior agreement. While these approaches are effective in structured domains such as Hanabi Bard et al. (2020) and Overcooked Carroll et al. (2019), they rely on symbolic observations, making them less suitable for complex, embodied human-agent collaboration. More recently, LLM-based agent frameworks Li et al. (2023); Yao et al. (2023); Zhang et al. (2024b; a); Liu et al. (2025) have demonstrated impressive capabilities in reasoning and planning within interactive environments. However, most assume known goals or static user preferences and lack mechanisms for inferring latent values through interaction. Our work builds on this line by addressing the challenge of fast adaptation to unknown, value-driven user goals via desire inference, memory utilization, and efficient, human-like dialogue in rich embodied tasks.
@@ 3 Embodied Home Assistance Simulation Environment
In real-world interactions, human users may provide vague instructions because their goals have only been partially formed and expressed at an abstract level Bettman et al. (1998). At the same time, people generally dislike reiterating their needs during communication Clark & Brennan (1991). Consequently, an effective assistant agent must be able to infer the latent goals and preferences of the human users and adapt its behavior accordingly, thereby minimizing redundant communication and execution costs to provide proactive assistance.
To study these challenges of agent–human adaptation in an embodied environment, we present HA-Desire (Home Assistance with diverse Desire), a novel embodied simulation environment designed to study agent-human adaptation in realistic home scenarios. As illustrated in Figure 2, HA-Desire builds upon VirtualHome Puig et al. (2018) and extends it with value-driven proxy human users. It includes 6 distinct home layouts containing typical rooms such as living rooms, kitchens, bathrooms, and bedrooms, with an average of 80 objects per room drawn from over 110 object classes. This diversity of layouts and object assets enables the construction of visually grounded tasks with high variability, providing a realistic testbed for embodied agent-human adaptation.
We instantiate a proxy human user within the environment. This user is assigned a set of latent value attributes and begins with only a vague task description. Guided by these values, the user samples a subset of desire-related goals from a larger goal space. Crucially, these goals are not explicitly revealed; instead, the proxy user provides only indirect hints about preferences and intentions in response to the ego agent’s inquiries.
@@
Figure 2: Overview of the HA-Desire environment. The simulation environment contains diverse objects and scenes. The proxy human user samples value attributes from a task-related space, which guides goal selection from a potential goal set via LLM. The user is constrained from directly revealing the true goals and instead communicates through desire-centered hints. This setup encourages the ego agent to infer user intent through interactive reasoning rather than relying on explicit instructions.
@@ 3.1 Problem Formulation
In the desire-centered agent-human adaptation problem, the task goal is not explicitly fixed, but the number of goals is predefined. Instead, there exists a potential goal set $G_{p}$, from which the task goal must be sampled by the human user $H$ and inferred by the ego agent $E$. Given a vague task description $T$, the proxy human user first samples a set of value attributes $V$ from a task-specific value space. Based on these values and the task description, the user then samples a set of desire-driven goals $G=G(V,G_{p},T)\subset G_{p}$.
@@ 3.2 Value-driven Human User
As illustrated in Figure 2, the proxy human user begins by sampling discretized value attributes $V$ from a task-related value space. For example, in the Prepare Snack task, the value space spans five dimensions: Hungry, Thirsty, SweetTooth, Fruitarian, and Alcoholic, each taking on one of three discrete levels—Not, Somewhat, or Very. These attributes reflect the user’s latent desires.
@@
To simulate natural and indirect human communication, the user is constrained via LLM prompting and output filtering to ensure that the true goal set is never revealed explicitly. Instead, it responds with value-driven hints that reflect its preferences and desires. This encourages the ego agent to reason about the user’s intent through interaction to provide proactive help with minimal repetitive annoying confirmation. The detailed LLM prompting strategy used to sample goals and generate user responses is provided in Appendix A.
Figure 3: Overview of the FAMER framework. FAMER comprises three key components: KeyInfo Extraction, Desire-Centered Mental Reasoning (including Goal Confirmation and Desire Inference), and Efficient Communication. These are supported by a memory module, perception module, and planning module, which together form an integrated pipeline for embodied agent-human adaptation.
4 Communication-Efficient Agent-Human Adaptation
To enable fast and communication-efficient adaptation to users, we propose a novel framework, FAMER (Fast Adaptation via MEntal Reasoning), which leverages the reasoning capabilities and commonsense embedded in large VLMs. In this work, we utilize GPT-4o Hurst et al. (2024).
As illustrated in Figure 3, FAMER integrates three core components: Key Information Extraction, Desire-Centered Mental Reasoning, and Efficient Communication. These modules work in concert to help the ego agent infer user intent, plan accurately, and minimize redundant interactions and communication. We detail each component in the following subsections.
@@ 4.1 Information Extraction
In HA-Desire, the ego agent receives first-person RGB-D images as observations. To extract structured information from these inputs, we employ a perception module based on Mask R-CNN He et al. (2017), trained on collected scene images following Zhang et al. (2024b). The module first predicts instance segmentation masks from the RGB image and then constructs 3D point clouds using the RGB-D data. These outputs are used to build a scene graph that encodes object locations, categories, and spatial relationships.
@@ 4.2 Desire-centered Mental Reasoning
The Goal Confirmation component extracts confirmed goals from the user’s responses by VLM reasoning. For example, if the agent asks, “Do you want some juice?” and the user replies, “Correct! Try to look for something crunchy,” the system confirms that juice is one of the user’s desired items. This process grounds part of the goal set and reduces future uncertainty.
@@ 4.3 Efficient Communication
When communication is necessary, the agent formulates targeted, desire-aligned questions aimed at resolving specific ambiguities. For example, if the agent has inferred a preference for sweet items but is uncertain about texture, it may ask “I found a cupcake. Would you like it?” instead of issuing vague or open-ended queries. This focused interaction strategy minimizes the communication burden on the user while enabling the agent to efficiently acquire high-value information.
@@ 5.1 Tasks & Metrics
We evaluate FAMER in two representative tasks instantiated within our proposed HA-Desire environment: Prepare Afternoon Snack and Set Up Dinner Table. Each task is associated with a five-dimensional value space that governs user preferences. The Snack task includes 10 potential goals, while the Table task contains 8. Each task is further divided into two levels: Medium and Large. In Medium setting, the number of target goals is 2, and the maximum episode length is 60 steps. In Large setting, the agent must satisfy 4 goals within 120 steps. For example, the Snack-M task involves a total of $C(10,2)=45$ possible goal combinations. The reason why we set the maximum goal space to 4 varying objects is that for daily tasks such as meal preparation, people often share a stable backbone of common objects and a small consideration subset of items ($\approx$3–5) that vary depending on individual preferences due to limited cognitive constraints Schank & Abelson (1975); Wood et al. (2022); Hauser & Wernerfelt (1990). Details of the tasks are provided in Appendix A.
For each task at a given level, every method is evaluated over 6 independent runs. In each run, the agent interacts with the human user for 3 episodes, with the user’s value attributes fixed throughout to match the agent-human adaptation setting. All reported results are averaged over the 6 runs.
@@
Score: Given $N$ total goals, the agent receives a reward of $\frac{1}{N}$ for each correct goal achieved. Completing an incorrect or distracting goal incurs a penalty of $\frac{1}{2N}$. The maximum achievable score per episode is 1, corresponding to the completion of all goals without any mistakes.
@@
Notably, HA-Desire supports diverse objects and scenes, and goal generation is automated, making it easy to extend to other tasks. Further details of the environments are provided in Appendix A.